Hello, I cannot figure out how to solve this geometric proof! Please help!
Prove that the length of the midline of a trapezoid is the average of the lengths of its two bases.

Answer by NancyLam(35)

length of the midline of a trapezoid is the sum of 2 base mutipfly 1/2
Example:The trapezoid QRST ,TS=12,QR=20 ,A and B are midle of the leg
AB=(12+20)/2 =32/2
AB= 16